I am thinking of three consecutive negative numbers. If i multiply the first with the second and then subtract three times the third, the result is 57. What are the numbers?
Answer by stanbon(75887)

I am thinking of three consecutive negative numbers. If i multiply the first with the second and then subtract three times the third, the result is 57. What are the numbers?
--------------------
1st: x-1
2nd: x
3rd: x+1
------------
Equations:
x(x-1)-3(x+1) = 57
x^2-x-3x-3 = 57
x^2-4x-60 = 0
(x-10)(x+6) = 0
Negative solution:
x = -6 (2nd)
x-1 = -7 (1st)
x+1 = -5 (3rd)
===================
Cheers,
Stan H.